I am supposed to solve by substitution. My equations are
3/4x + 1/2y = -7/12
and
x - y = -4/3
I have tried getting x alone on the first question but can't get the right answer

Answer by TimothyLamb(4379)

---
3/4x + 1/2y = -7/12
x - y = -4/3
---
put the system of linear equations into standard form:
---
( 3/4x + 1/2y = -7/12 ) 12
( x - y = -4/3 ) 3
---
9x + 6y = -7
3x - 3y = -4
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = -1
y = 0.33333333 = 1/3
---
system matrix
9 6
3 -3
---
inverse of system matrix
0.066666667 0.13333333
0.066666667 -0.2
---
determinant of system matrix = -45
---
now solve by substitution method:
---
6y = -7 - 9x
y = -(7/6) - (9/6)x
---
substitute into:
3x - 3y = -4
3x - 3( -(7/6) - (9/6)x ) = -4
3x + (7/2) + (9/2)x = -4
3x + (9/2)x = -4 - (7/2)
(6/2)x + (9/2)x = -(8/2) - (7/2)
( (6/2)x + (9/2)x = -(8/2) - (7/2) ) 2
15x = -15
---
x = -1
---
y = -(7/6) - (9/6)x
y = -(7/6) - (9/6)(-1)
y = -(7/6) + (9/6)
y = 2/6
---
y = 1/3
---
solution by substitution method:
x = -1
y = 0.33333333 = 1/3
---
both methods agree!
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php